# Multi-modal Generative AI: Multi-modal LLMs, Diffusions and the Unification

### v2

@@ IV-B1 Multi-modal input processor
To tackle the multi-modal input text and images, two possible input processors are presented in Fig. 10. Text is consistently tackled by a text tokenizer. However, there are some differences in the visual input. In Fig. 10(a), we show the visual processor adopted by most existing works, where a single visual encoder is used to process the images. Considering that the visual tokens should support the pixel-level visual generation task, existing works [3, 152, 151] generally adopt the single pixel-level (or patch-level) visual tokens. The pixel-level tokens bring challenges to the multi-modal transformer, requiring it not only to capture the relations between image patches for visual generation but also to visual abstract reasoning ability for understanding. In contrast, a possible alternative multi-modal input processor is presented in Fig. 10(b). For each image, we respectively use a semantic encoder and a pixel-level encoder to obtain both semantic and pixel tokens. For the generation objective, the visual-pixel tokens are used. While optimizing the understanding objective, the visual semantic tokens are activated. Janus-pro [154] and BAGEL [155] adopt this architecture and outperform existing unified works. Moreover, it is a more flexible way to conduct some adaptive token selection from the semantic and pixel tokens for fine-grained understanding. We believe this would result in interesting research work.
@@ IV-B2 Multi-modal Transformer
After discussing how to tackle the multi-modal input information, the next key component is the multi-modal transformer, which captures the complex relations among and within modalities. As shown in Fig. 11, on the left is a dense model, where one unified transformer is used for both multi-modal understanding and generation [156, 41]. Considering that understanding and generation might share some knowledge but their objectives are not exactly the same, it is a natural idea to utilize the mixture of experts [157] in multi-task learning as shown in (b). On the right of the figure, some of the experts share the knowledge of understanding and generation, e.g., concepts and their relations, some of the experts are good at the visual details for visual generation, and other experts are good at reasoning for better understanding. LlamaFusion [158] and BAGEL [155] have made preliminary explorations in this area, both using only two experts and employing hard routing. In LlamaFusion, which uses a single visual encoder, one expert is responsible for processing text tokens, while the other handles visual tokens. In contrast, BAGEL, which adopts semantic-pixel visual encoders, assigns one expert to process text tokens and visual semantic tokens, and the other to handle visual pixel tokens. Both works find that their architectures outperform dense models, indicating that unified models still face optimization challenges arising from task-specific or modality-specific learning objectives.
@@
In this section, we provide a discussion of the unified model of multi-modal generation and multi-modal understanding, from both the probabilistic modeling methods and the model architectures. The discussed techniques can also be combined with each other and result in more architectures, and now there are few attempts at the unified model design and we believe the discussion above will inspire a lot of future works.
@@
Model Type Classification Publication Year
@@
LLaVA [4] Image LLM Alignment NeurIPS 2024
BLIP-2 [24] Image LLM Alignment ICML 2023
MiniGPT-4 [30] Image LLM Alignment ICLR 2024
Qwen-VL [159] Image LLM Alignment ArXiv 2023
Flamingo [29] Image LLM Alignment NeurIPS 2025
Fuyu [33] Image LLM Early-Fusion - 2023
Gemini [25] Image LLM Early-Fusion ArXiv 2023
VideoChat [48] Video LLM Alignment ArXiv 2023
VideoLLaMA [49] Video LLM Alignment EMNLP 2023
Video-ChatGPT [50] Video LLM Alignment ACL 2023
VideoLLaMA2 [51] Video LLM Alignment ArXiv 2024
LLaVA-OneVision [15] Video LLM Alignment TMLR 2024
MiniCPM-V [160] Video LLM Alignment ArXiv 2024
VILA-1.5 [161] Video LLM Alignment ArXiv 2023
@@
GLIDE [72] Text-to-Image Pixel-Based ICML 2022
Imagen [73] Text-to-Image Pixel-Based NeurIPS 2022
Stable Diffusion [74] Text-to-Image Latent-Based CVPR 2022
DALL-E2 [75] Text-to-Image Latent-Based ArXiv 2022
DiT [78] Text-to-Image Latent-Based ICCV 2023
PixArt-$\alpha$ [79] Text-to-Image Latent-Based ICLR 2025
Text2Video-Zero [103] Text-to-Video Latent-Based ICCV 2023
Latent-Shift [104] Text-to-Video Latent-Based ArXiv 2023
VDM [105] Text-to-Video Latent-Based NeurIPS 2022
Make-a-video [106] Text-to-Video Latent-Based ICLR 2024
AnimateDiff [107] Text-to-Video Latent-Based ICLR 2024
Latte [108] Text-to-Video Latent-Based TMLR 2025
CogVideo [162] Text-to-Video Latent-Based ICLR 2023
Vidu [133] Text-to-Video Latent-Based ArXiv 2024
@@
VL-GPT [136] Unified Model Autoregressive ArXiv 2023
Chameleon [3] Unified Model Autoregressive ArXiv 2024
Emu3 [44] Unified Model Autoregressive ArXiv 2024
LlamaGen [137] Unified Model Autoregressive ArXiv 2024
AnyGPT [138] Unified Model Autoregressive ACL 2024
Janus  [139] Unified Model Autoregressive CVPR 2025
Janus-Pro  [154] Unified Model Autoregressive ArXiv 2025
VisualGPT [143] Unified Model Joint AR-Diffusion ArXiv 2023
HuggingGPT [144] Unified Model Joint AR-Diffusion NeurIPS 2024
MLLM-Tool [145] Unified Model Joint AR-Diffusion WACV 2025
Emu2 [146] Unified Model Joint AR-Diffusion CVPR 2024
Kosmos-G [147] Unified Model Joint AR-Diffusion ICLR 2024
CoDi-2 [148] Unified Model Joint AR-Diffusion CVPR 2024
Seed-X [149] Unified Model Joint AR-Diffusion ArXiv 2024
MetaQuery [163] Unified Model Joint AR-Diffusion ArXiv 2025
BLIP3-o [150] Unified Model Joint AR-Diffusion ArXiv 2025
TransFusion [151] Unified Model Joint AR-Diffusion ICLR 2025
Show-o [152] Unified Model Joint AR-Diffusion ICLR 2025
LlamaFusion [158] Unified Model Joint AR-Diffusion Arxiv 2024
BAGEL [155] Unified Model Joint AR-Diffusion Arxiv 2025
@@ V Datasets
TABLE II: Common datasets
@@ V-A Caption Datasets
The caption dataset aims to improve basic visual and temporal description capabilities for multi-modal LLMs and provide the mapping relationship for text-to-image and text-to-video models. Commonly used text-to-image datasets include SBU Captions [165], MSCOCO [166], Conceptual Captions (CC-3M) [167], and LAION [168]. The size of these datasets ranges from 328K to 5B. Recently MINT-1T is proposed comprising one trillion text tokens and three billion images [169], a 10x scale-up from existing open-source datasets and it includes previously untapped sources such as PDFs and ArXiv papers. Text-to-video datasets include WebVid [170], InternVid [171], HD-VG-130M [172], YouCook2 [173], and TextVR [174].
The caption datasets mainly serve in the following two aspects, i.e., (i) provide knowledge for the training of generation models to generate images or videos based on the input text embedding and (ii) use text-image datasets to align the image modality with the multi-modal LLM for understanding visual inputs.
@@ V-B Conversation Datasets
The conversation dataset aims at enhancing multi-modal LLMs’ capabilities for single-turn and multi-turn conversations when asking questions about the input image or video. Normally a diverse set of questions would be asked about the visual content of the image and the video, including the object types, counting the objects, objects actions, objects locations, event moment, event duration, and relative positions between objects. With simple formatting reorganization, many visual QA datasets could be directly constructed as conversation datasets for multi-modal LLM training. These include basic VQA (VQAv2 [175], GQA [176]), knowledge-based VQA (OK-VQA [177], AOK-VQA [178]), OCR-based VQA (OCR-VQA [179], TextVQA [180]) and VideoQA (TGIF-QA [181], WebVidQA [182], and egocentric VQA from Ego4D [183]), which can not only improve the visual QA capabilities for multi-modal LLMs in conversations but also help the models to learn more visual and temporal knowledge.
@@ V-D Integration Datasets
Due to the strong generalization ability of multi-modal LLMs, their training data is not limited to only one single task such as caption, conversation, or reasoning, instead requiring comprehensive pretraining for both simple and complex visual modal tasks. Therefore, many multi-modal large model works often do not use a single visual task dataset, instead they select subsets of several datasets from each category mentioned above for integration and adjustment, forming instruction training data that employs both image and video data for different visual modal tasks. For visual instruction tuning, LLaVA [28] is the first multi-modal LLM which i) leverages text-only GPT-4 [189] to expand the existing bounding box, and ii) employs caption dataset (e.g., MSCOCO [166]) as multi-modal instruction tuning data. In addition, Liu et al. propose LLaVA-Instruct, which is built on a subset of the CC-3M dataset and contains 58k in conversations, 23k in detailed descriptions, as well as 77k in complex reasoning records. Following the development of visual instruction tuning, many video LLMs such as Video-LLaVA [156], VideoChat2 [188] and VideoLLaMa2 [51] are proposed, utilizing the combination of caption, conversation, and reasoning datasets under both text-image and text-video modalities.
@@ VI-A Unified Model for Video Understanding and Generation
In Section IV, we primarily discuss the unified models for image understanding and generation. Given the large amount of video data in the wild, we believe there will be an urging need to extend the unification to videos. Among the three architectures introduced in Fig. 9, bridging the multi-modal LLM and video diffusion model with a connector [190, 191] can be achieved in a way similar to images. However, adapting the other two architectures to videos faces significant challenges due to i) the increased computational demands caused by longer sequences as well as ii) the difficulty in learning spatiotemporal cues. For instance, in an autoregressive model, encoding individual video frames separately using a 2D visual tokenizer fails to capture the essential temporal motion information. VideoPoet [192], which employs a 3D video tokenizer [193], encodes a 17-frame video (spanning 2.125 seconds) into 1280 tokens, limiting its ability to generate longer videos. VideoLaViT [194] introduces an efficient video representation model by decomposing videos into keyframes and temporal motions, training separate tokenizers for each of them, which significantly improves computational efficiency. However, the training cost is still too high when scaling to the large amount of web-scale video data. Similarly, using a single model trained with both diffusion and autoregressive regularizations also encounters the same challenges, where modeling complex relations such as causal attention and spatiotemporal attention within the model remains unexplored. Therefore, it deserves more efforts in advancing unified generative AI for video understanding and generation.

### v3

@@ IV-B1 Multi-modal input processor
To tackle the multi-modal input text and images, two possible input processors are presented in Fig. 9. Text is consistently tackled by a text tokenizer. However, there are some differences in the visual input. In Fig. 9(a), we show the visual processor adopted by most early works, where a single visual encoder is used to process the images. Considering that the visual tokens should support the pixel-level visual generation task, early works [3, 134, 133] generally adopt the single pixel-level (or patch-level) visual tokens (e.g., VQVAE). The pixel-level tokens bring challenges to the multi-modal transformer, requiring it not only to capture the relations between image patches for visual generation but also to visual abstract reasoning ability for understanding. In contrast, a possible alternative multi-modal input processor is presented in Fig. 9(b). For each image, we respectively use a semantic encoder (e.g., CLIP-ViT) and a pixel-level encoder (e.g., VQVAE) to obtain both semantic and pixel tokens. Janus [118] was the first to adopt this architecture. It introduced two separate visual encoding paths: a semantic encoder for extracting visual features in understanding tasks, and a pixel-level encoder for encoding images in generation tasks. Subsequent works, such as UniToken [136], further explored directly concatenating features of the two encoders along the sequence dimension, allowing the model to receive both types of features simultaneously for understanding and generation tasks. By using a dual-encoder approach, models can leverage both low-level pixel information and high-level semantic information, which better enhances performance on both understanding and generation tasks. Consequently, most recent works adopt this architecture. Moreover, it is a more flexible way to conduct some adaptive token selection from the semantic and pixel tokens for fine-grained understanding. We believe this would result in interesting research work.
@@ IV-B2 Multi-modal Transformer
After discussing how to tackle the multi-modal input information, the next key component is the multi-modal transformer, which captures the complex relations among and within modalities. As shown in Fig. 10, on the left is a dense model, where one unified transformer is used for both multi-modal understanding and generation [137, 44]. Considering that understanding and generation might share some knowledge but their objectives are not exactly the same, it is a natural idea to utilize the mixture of experts [138] in multi-task learning as shown in (b). On the right of the figure, some of the experts share the knowledge of understanding and generation, e.g., concepts and their relations, some of the experts are good at analyzing visual details for visual generation, and other experts are good at conducting reasoning for better understanding. LlamaFusion [139] and BAGEL [124] have made preliminary explorations in this area, both using only two experts and employing hard routing. In LlamaFusion, which uses a single visual encoder, one expert is responsible for processing text tokens, while the other handles visual tokens. In contrast, BAGEL, which adopts semantic-pixel visual encoders, assigns one expert to process text tokens and visual semantic tokens, and the other to handle visual pixel tokens. Both works find that their architectures outperform dense models, indicating that unified models still face optimization challenges arising from task-specific or modality-specific learning objectives.
@@
In Table II, we present the performance of several recent unified models. Due to large differences in model size and training data volume, a fair comparison is difficult. Regarding architecture choice: currently, there are still no large-scale Autoregressive Models trained with massive data. The latest Skywork UniPic demonstrates strong capabilities in generation and editing, but its performance on understanding tasks is not reported. In the Connector-based Joint Models category, MetaQueries, BLIP3o, and Qwen-Image all adopt Qwen2.5-vl-7B as the MLLM, resulting in similar performance on understanding tasks. However, the success of Qwen-Image indicates that increasing the scale of the diffusion model and enlarging the training dataset can significantly boost performance in generation and editing tasks. In the Autoregressive-Diffusion Joint Models category, BAGEL leverages the largest model and dataset, making it a strong competitor to Qwen-Image. Regarding the choice of visual encoder: most recent models adopt the dual encoder (Semantic-Pixel Visual Encoders) architecture, which benefits both understanding and generation tasks. Since models using MoE are still limited, it remains unclear whether MoE brings significant advantages. We hope that future work will explore this direction further.
In this section, we provide a discussion of the unified model of multi-modal generation and multi-modal understanding, from both the probabilistic modeling methods and the model architectures. Though the discussed techniques can combine with each other to form more architectures as well, there are very few attempts at the unified model design, making us believe the inspirations of many future works brought by the discussions above.
@@
Model Institution Type Classification Publication Year Parameters
@@
LLaVA [4] Microsoft Image LLM Alignment NeurIPS 2024 13B
BLIP-2 [28] Salesforce Image LLM Alignment ICML 2023 12B
MiniGPT-4 [33] KAUST Image LLM Alignment ICLR 2024 7B
Qwen-VL [140] Alibaba Image LLM Alignment ArXiv 2023 7B
Flamingo [32] DeepMind Image LLM Alignment NeurIPS 2025 3B
Fuyu [36] Adept Image LLM Early-Fusion - 2023 8B
Gemini [29] Google Image LLM Early-Fusion ArXiv 2023 -
Claude3 [141] Anthropic Image LLM Early-Fusion - 2024 -
VideoChat [51] Shanghai AI Lab Video LLM Alignment ArXiv 2023 7B
VideoLLaMA [52] Alibaba Video LLM Alignment EMNLP 2023 7B
VideoLLaMA2 [54] Alibaba Video LLM Alignment ArXiv 2024 7B
Video-ChatGPT [53] MBZUAI Video LLM Alignment ACL 2023 7B
LLaVA-OneVision [19] ByteDance Video LLM Alignment TMLR 2024 7B
MiniCPM-V [142] OpenBMB Video LLM Alignment ArXiv 2024 8B
VILA-1.5 [143] NVIDIA Video LLM Alignment ArXiv 2023 7B
Pengi [144] Microsoft Speech LLM Alignment NeurIPS 2023 1B
Salmonn [145] ByteDance Speech LLM Alignment ICLR 2024 13B
Qwen-Audio [146] Alibaba Speech LLM Alignment ArXiv 2023 7B
OSUM [147] NPU Speech LLM Alignment ArXiv 2025 7B
VALL-E [71] Microsoft Speech LLM Early-Fusion ArXiv 2025 7B
SpeechGPT [72] Fudan University Speech LLM Early-Fusion EMNLP 2023 7B
AudioPaLM [73] Google Speech LLM Early-Fusion ArXiv 2023 8B
@@
GLIDE [87] OpenAI Text-to-Image Pixel-Based ICML 2022 5B
Imagen [88] Google Text-to-Image Pixel-Based NeurIPS 2022 3B
Stable Diffusion [27] LMU Text-to-Image Latent-Based CVPR 2022 1B
DALL-E2 [89] OpenAI Text-to-Image Latent-Based ArXiv 2022 6B
DiT [95] Meta Text-to-Image Latent-Based ICCV 2023 1B
PixArt-$\alpha$ [96] Huawei Text-to-Image Latent-Based ICLR 2025 1B
Text2Video-Zero [99] Picsart AI Text-to-Video Latent-Based ICCV 2023 1B
Latent-Shift [100] Meta Text-to-Video Latent-Based ArXiv 2023 2B
VDM [101] Google Text-to-Video Latent-Based NeurIPS 2022 -
Make-a-video [102] Meta Text-to-Video Latent-Based ICLR 2024 10B
AnimateDiff [103] Shanghai AI Lab Text-to-Video Latent-Based ICLR 2024 1B
Latte [104] Shanghai AI Lab Text-to-Video Latent-Based TMLR 2025 1B
CogVideo [148] Z.AI Text-to-Video Latent-Based ICLR 2023 15B
Wan [149] Alibaba Text-to-Video Latent-Based ArXiv 2025 14B
HunyuanVideo [150] Tencent Text-to-Video Latent-Based ArVix 2024 13B
Vidu [151] Shengshu Text-to-Video Latent-Based ArXiv 2024 -
DiffWave [107] Baidu Text-to-Speech Vocoder ICLR 2021 6M
WaveGrad [108, 111] Google Text-to-Speech Vocoder ICLR 2021 23M
Grad-TTS [109] Huawei Text-to-Speech Acoustic Model ICML 2021 30M
Diff-TTS [110] Neosapience Text-to-Speech Acoustic Model Interspeech 2021 13M
@@
VL-GPT [114] Tencent Unified Model Autoregressive ArXiv 2023 8B
Chameleon [3] Meta Unified Model Autoregressive ArXiv 2024 7B
Emu2 [117] BAAI Unified Model Autoregressive CVPR 2024 37B
Emu3 [47] BAAI Unified Model Autoregressive ArXiv 2024 8B
LlamaGen [115] ByteDance Unified Model Autoregressive ArXiv 2024 3B
AnyGPT [116] Shanghai AI Lab Unified Model Autoregressive ACL 2024 8B
Janus [118] DeepSeek Unified Model Autoregressive CVPR 2025 1B
Janus-Pro [122] DeepSeek Unified Model Autoregressive ArXiv 2025 7B
Skywork UniPic [152] Skywork Unified Model Autoregressive ArXiv 2025 2B
VisualGPT [125] Microsoft Unified Model Joint AR-Diffusion ArXiv 2023 -
HuggingGPT [126] Microsoft Unified Model Joint AR-Diffusion NeurIPS 2024 -
MLLM-Tool [127] Meituan Unified Model Joint AR-Diffusion WACV 2025 13B
Kosmos-G [128] Microsoft Unified Model Joint AR-Diffusion ICLR 2024 2B
CoDi-2 [129] Microsoft Unified Model Joint AR-Diffusion CVPR 2024 8B
Seed-X [130] Tencent Unified Model Joint AR-Diffusion ArXiv 2024 13B
MetaQuery [153] Meta Unified Model Joint AR-Diffusion ArXiv 2025 7B
BLIP3o [131] Salesforce Unified Model Joint AR-Diffusion ArXiv 2025 8B
OmniGen2 [154] BAAI Unified Model Joint AR-Diffusion ArXiv 2025 7B
Qwen-Omni [155, 156] Alibaba Unified Model Joint AR-Diffusion ArXiv 2025 30B
Ming-Omni [157] Ant Group Unified Model Joint AR-Diffusion ArXiv 2025 7B
TransFusion [133] Meta Unified Model Joint AR-Diffusion ICLR 2025 7B
Show-o [134] NUS Unified Model Joint AR-Diffusion ICLR 2025 1B
Show-o2 [158] NUS Unified Model Joint AR-Diffusion ArXiv 2025 7B
LlamaFusion [139] Meta Unified Model Joint AR-Diffusion Arxiv 2024 8B
Mogao [123] ByteDance Unified Model Joint AR-Diffusion Arxiv 2025 7B
BAGEL [124] ByteDance Unified Model Joint AR-Diffusion Arxiv 2025 7B
TABLE II: Comparison of recent multi-modal models across understanding, generation, and editing benchmarks.
Model Date Params Data Dual Encoder MoE Understanding Generation Editing
MMBench MMMU $M$-Vet WISE GenEval DPGBench ImgEdit GEdit-Bench-EN
[…]
Autoregressive Models
Emu3 [47] 2024.9 8B - $\times$ $\times$ 58.5 31.6 37.2 0.39 0.66 80.6 - -
Janus-Pro [122] 2025.1 7B 144M $\times$ 79.2 41.0 50.0 0.35 0.80 84.19 - -
Skywork UniPic [152] 2025.8 2B 130M $\times$ - - - - 0.86 85.50 3.49 5.83
Connector-based Joint Models
MetaQueries [153] 2025.4 7B+1.6B 25M $\times$ 83.5 58.6 66.6 0.55 0.80 82.05 - -
BLIP3o [131] 2025.5 7B+1.4B 25M $\times$ 83.5 50.6 66.6 0.62 0.84 81.6 - -
OmniGen2 [154] 2025.6 3B+4B 66M $\times$ 79.1 53.1 61.8 - 0.80 83.57 3.44 6.42
Qwen-Image [132] 2025.8 7B+20B >>1000M $\times$ 83.5 58.6 67.1 - 0.87 88.32 4.27 7.56
Autoregressive-Diffusion Joint Models
Mogao [123] 2025.5 7B - 75.0 44.2 - - 0.89 84.33 - -
BAGEL [124] 2025.5 14B 1600M 85.0 55.3 67.2 0.52 0.88 85.07 3.20 6.52
Show-o2 [158] 2025.6 7B 66M $\times$ 79.3 48.9 - - 0.76 86.14 - -
@@ V Datasets
TABLE III: Common datasets
@@ V-A Caption Datasets
The caption dataset aims to improve basic visual and temporal description capabilities for multi-modal LLMs and provide the mapping relationship for text-to-image and text-to-video models. Commonly used text-to-image datasets include SBU Captions [160], MSCOCO [161], Conceptual Captions (CC-3M) [162], and LAION [163]. The size of these datasets ranges from 328K to 5B. Recently, MINT-1T has been proposed, comprising one trillion text tokens and three billion images [164], a 10x scale-up from existing open-source datasets, and it includes previously untapped sources such as PDFs and ArXiv papers. Text-to-video datasets include WebVid [165], InternVid [166], HD-VG-130M [167], YouCook2 [168], and TextVR [169].
The caption datasets mainly serve in the following two aspects, i.e., (i) provide knowledge for the training of generation models to generate images or videos based on the input text embedding, and (ii) use text-image datasets to align the image modality with the multi-modal LLM for understanding inputs.
@@ V-B Conversation Datasets
The conversation dataset aims at enhancing multi-modal LLMs’ capabilities for single-turn and multi-turn conversations when asking questions about the input image or video. Normally, a diverse set of questions would be asked about the visual content of the image and the video, including the object types, counting the objects, object actions, object locations, event moment, event duration, and relative positions between objects. With simple formatting reorganization, many visual QA datasets could be directly constructed as conversation datasets for multi-modal LLM training. These include basic VQA (VQAv2 [48], GQA [170]), knowledge-based VQA (OK-VQA [171], AOK-VQA [172]), OCR-based VQA (OCR-VQA [173], TextVQA [174]) and VideoQA (TGIF-QA [175], WebVidQA [176], and egocentric VQA from Ego4D [177]), which can not only improve the visual QA capabilities for multi-modal LLMs in conversations but also help the models to learn more visual and temporal knowledge.
@@ V-D Integration Datasets
Due to the strong generalization ability of multi-modal LLMs, their training data is not limited to only one single task, such as caption, conversation, or reasoning, instead requiring comprehensive pretraining for both simple and complex visual modal tasks. Therefore, many multi-modal large model works often do not use a single visual task dataset. Instead, they select subsets of several datasets from each category mentioned above for integration and adjustment, forming instruction training data that employs both image and video data for different visual modal tasks. For visual instruction tuning, LLaVA [4] is the first multi-modal LLM, which i) leverages text-only GPT-4 [1] to expand the existing bounding box, and ii) employs caption dataset (e.g., MSCOCO [161]) as multi-modal instruction tuning data. In addition, Liu et al. propose LLaVA-Instruct, which is built on a subset of the CC-3M dataset and contains 58k in conversations, 23k in detailed descriptions, as well as 77k in complex reasoning records. Following the development of visual instruction tuning, many video LLMs such as Video-LLaVA [137], VideoChat2 [182], and VideoLLaMa2 [54], are proposed, utilizing the combination of caption, conversation, and reasoning datasets under both text-image and text-video modalities.
@@ VI-A Unified Model for Video Understanding and Generation
In Section IV, we primarily discuss the unified models for image understanding and generation. Given the large amount of video data in the wild, we believe there will be an urgent need to extend the unification to videos [183, 184, 185]. Among the three architectures introduced in Fig. 8, bridging the multi-modal LLM and video diffusion model with a connector [186, 187] can be achieved in a way similar to images. However, adapting the other two architectures to videos faces significant challenges due to i) the increased computational demands caused by longer sequences, as well as ii) the difficulty in learning spatiotemporal cues. For instance, in an autoregressive model, encoding individual video frames separately using a 2D visual tokenizer fails to capture the essential temporal motion information. VideoPoet [188], which employs a 3D video tokenizer [189], encodes a 17-frame video (spanning 2.125 seconds) into 1280 tokens, limiting its ability to generate longer videos. VideoLaViT [190] introduces an efficient video representation model by decomposing videos into keyframes and temporal motions, training separate tokenizers for each of them, which significantly improves computational efficiency. However, the training cost is still too high when scaling to the large amount of web-scale video data. Similarly, using a single model trained with both diffusion and autoregressive regularizations also encounters the same challenges, where modeling complex relations such as causal attention and spatiotemporal attention within the model remains unexplored. Therefore, it deserves more effort in advancing unified generative AI for video understanding and generation.